# Equivalence Relations on Vertex Operator Algebras, II: Witt Equivalence and Orbifolds

### v1

@@ 7.4. Examples
Because $U$ and $W$ contain $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}}$ as a subalgebra, we may describe them by decomposing them into $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}}$-modules by restriction. For example, $U$ decomposes into $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}}$-modules as
$U\cong\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}\oplus\mathcal{P}_{3}(2,0)\otimes\textsl{VF}_{24}^{\natural}(2,0),$
where we have used the notation of [BHL+21] to label the modules of $\mathcal{P}_{3}$ and $\smash{\textsl{VF}_{24}^{\natural}}$. Let us do the same for $W$ in more detail. We may obtain it and its irreducible modules as simple-current extensions of $\smash{\mathcal{P}_{3}\otimes\textsl{VF}_{24}^{\natural}}$,
(9) $\begin{split}W_{(\mathbf{1},\mathbf{1})}&=\mathcal{P}_{3}\textsl{VF}_{24}^{\natural}\oplus\mathcal{P}_{3}(3,1)\textsl{VF}_{24}^{\natural}(3,1)\oplus\mathcal{P}_{3}(3,-1)\textsl{VF}_{24}^{\natural}(3,-1),\\ W_{(\mathbf{1},\tilde{\tau})}&=\mathcal{P}_{3}\textsl{VF}_{24}^{\natural}(2,0)\oplus\mathcal{P}_{3}(3,1)\textsl{VF}_{24}^{\natural}(1,1)\oplus\mathcal{P}_{3}(3,-1)\textsl{VF}_{24}^{\natural}(2,2),\\ W_{(\tau,\mathbf{1})}&=\mathcal{P}_{3}(2,0)\textsl{VF}_{24}^{\natural}\oplus\mathcal{P}_{3}(1,1)\textsl{VF}_{24}^{\natural}(3,1)\oplus\mathcal{P}_{3}(2,2)\textsl{VF}_{24}^{\natural}(3,-1),\\ W_{(\tau,\tilde{\tau})}&=\mathcal{P}_{3}(2,0)\textsl{VF}_{24}^{\natural}(2,0)\oplus\mathcal{P}_{3}(1,1)\textsl{VF}_{24}^{\natural}(1,1)\oplus\mathcal{P}_{3}(2,2)\textsl{VF}_{24}^{\natural}(2,2),\end{split}$
where to avoid clutter we have suppressed the symbol $\otimes$. Here, we are labeling the modules of $W$ by the simple objects $(\mathbf{1},\mathbf{1})$, $(\mathbf{1},\tilde{\tau})$, $(\tau,\mathbf{1})$ and $(\tau,\tilde{\tau})$ of $\operatorname{Rep}(W)\cong Z(\operatorname{Fib})\cong(G_{2},1)\boxtimes(F_{4},1)$, where $\tau$ and $\tilde{\tau}$ are the non-trivial simple objects of $(G_{2},1)$ and $(F_{4},1)$, respectively. One may indeed check that the character vector of this $W$ matches equation (7) by using equations (3.20) and (3.25) of [BHL+21]. [^19] We note a typo in the transcription of the parafermion characters in version $2$ of [BHL+21] on the arXiv and in the published version. The correct characters are $\psi^{(3)}_{1,1}(\tau)=\psi^{(3)}_{2,2}(\tau)=q^{1/30}(1+q+2q^{2}+3q^{3}+5q^{4}+7q^{5}+\dots)$ and $\psi_{2,0}^{(3)}(\tau)=q^{11/30}(1+2q+2q^{2}+4q^{3}+5q^{4}+8q^{5}+\dots)$.
The characters of $V^{\natural}$ twisted by the Fibonacci symmetry corresponding to $W$ can be computed in terms of the ordinary characters of $W$, which can in turn be computed in terms of the characters of $\mathcal{P}_{3}$ and $\textsl{VF}_{24}^{\natural}$ using the identifications in equation (9). For example, using $\smash{\hat{\mathcal{L}}}$ to denote the operator acting on $\smash{V^{\natural}}$ corresponding to the non-trivial line $\mathcal{L}$ in the Fibonacci category $\operatorname{Fib}$, one obtains
@@
where $\varphi=\frac{1+\sqrt{5}}{2}$. One may then compute the $\mathcal{L}$-twisted sector $\smash{V^{\natural}_{\mathcal{L}}}$ by taking a modular $S$-transformation of equation (10). Alternatively, one may proceed abstractly by noting that $\smash{V^{\natural}}$ defines a Lagrangian algebra $L$ inside $\operatorname{Rep}(W)$, and the category $\smash{\operatorname{Ver}(V^{\natural}/W)\cong\operatorname{Fib}}$ is given by $\operatorname{Rep}(W)_{L}$, the category of $L$-modules inside $\operatorname{Rep}(W)$. It is known that there are precisely two irreducible $L$-modules, $L$ and $M$, in $\operatorname{Rep}(W)$, corresponding to the trivial line and the Fibonacci line $\mathcal{L}$, respectively. They decompose as simple objects of $\operatorname{Rep}(W)$ as
@@
It would be interesting to identify a Fibonacci symmetry of $V^{\natural}$ that corresponds to a subalgebra realizing the second option, equation (8), as its character vector.
@@
Let $\tilde{\alpha}=\frac{1}{\sqrt{2k}}$. Following [DN99b], the irreducible modules of $V^{+}_{\sqrt{k}A_{1}}$ are
$\bigl{\{}V_{\sqrt{k}A_{1}}^{\pm},V_{\sqrt{k}A_{1}+k\tilde{\alpha}}^{\pm}\bigr{\}}\cup\bigl{\{}V_{\sqrt{k}A_{1}+i\tilde{\alpha}}\mid i=1,\dots,k-1\bigr{\}}\cup\bigl{\{}V_{\sqrt{k}A_{1}}^{T_{i},\pm}\mid i=1,2\bigr{\}},$
and it is straightforward to determine that $V_{A_{1}}$ and its non-trivial irreducible module $V_{A_{1}}(1/4)$ decompose as follows. If $\ell$ is even, then
@@
$V_{A_{1}}(1/4)$ $\cong\bigoplus_{j=0}^{\ell/2-1}2\cdot V_{\ell A_{1}+(2j+1)\ell\tilde{\alpha}},$
@@
$V_{A_{1}}(1/4)$ $\cong V_{\ell A_{1}+\ell^{2}\tilde{\alpha}}^{+}\oplus V_{\ell A_{1}+\ell^{2}\tilde{\alpha}}^{-}\oplus\bigoplus_{j=0}^{(\ell-3)/2}2\cdot V_{\ell A_{1}+(2j+1)\ell\tilde{\alpha}}.$
@@
$\mathcal{H}_{\mathrm{SU}(2)_{1}}\cong V_{A_{1}}\otimes\overline{V_{A_{1}}}\oplus V_{A_{1}}(1/4)\otimes\overline{V_{A_{1}}(1/4)}$

### v2

@@ 7.4. Examples
Because $U$ and $W$ contain $\smash{\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural}}$ as a conformal subalgebra, we may decompose them into modules for it. For example, $U$ decomposes as
$U\cong\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural}\oplus\mathcal{P}(3,[2,0])\otimes\textsl{VF}_{24}^{\natural}(2,0),$
where we have used the notation of [BHL+21] to label the modules of $\mathcal{P}(3)$ and $\smash{\textsl{VF}_{24}^{\natural}}$. Let us do the same for $W$ in more detail. We may obtain it and its irreducible modules as simple-current extensions of $\smash{\mathcal{P}(3)\otimes\textsl{VF}_{24}^{\natural}}$,
(9) $\begin{split}W_{(\mathbf{1},\mathbf{1})}&=\mathcal{P}(3)\textsl{VF}_{24}^{\natural}\oplus\mathcal{P}(3,[3,1])\textsl{VF}_{24}^{\natural}(3,1)\oplus\mathcal{P}(3,[3,-1])\textsl{VF}_{24}^{\natural}(3,-1),\\ W_{(\mathbf{1},\tilde{\tau})}&=\mathcal{P}(3)\textsl{VF}_{24}^{\natural}(2,0)\oplus\mathcal{P}(3,[3,1])\textsl{VF}_{24}^{\natural}(1,1)\oplus\mathcal{P}(3,[3,-1])\textsl{VF}_{24}^{\natural}(2,2),\\ W_{(\tau,\mathbf{1})}&=\mathcal{P}(3,[2,0])\textsl{VF}_{24}^{\natural}\oplus\mathcal{P}(3,[1,1])\textsl{VF}_{24}^{\natural}(3,1)\oplus\mathcal{P}(3,[2,2])\textsl{VF}_{24}^{\natural}(3,-1),\\ W_{(\tau,\tilde{\tau})}&=\mathcal{P}(3,[2,0])\textsl{VF}_{24}^{\natural}(2,0)\!\oplus\!\mathcal{P}(3,[1,1])\textsl{VF}_{24}^{\natural}(1,1)\!\oplus\!\mathcal{P}(3,[2,2])\textsl{VF}_{24}^{\natural}(2,2),\end{split}$
where to avoid clutter we have suppressed the symbol $\otimes$. Here, we are labeling the irreducible modules of $W=W_{(\mathbf{1},\mathbf{1})}$ by the simple objects $(\mathbf{1},\mathbf{1})$, $(\mathbf{1},\tilde{\tau})$, $(\tau,\mathbf{1})$ and $(\tau,\tilde{\tau})$ of $\operatorname{Rep}(W)\cong Z(\mathrm{Fib})\cong(G_{2},1)\boxtimes(F_{4},1)$, where $\tau$ and $\tilde{\tau}$ are the non-trivial simple objects of $(G_{2},1)$ and $(F_{4},1)$, respectively. One may indeed check that the character vector of this $W$ matches equation (7) by using equations (3.20) and (3.25) of [BHL+21]. [^20] We note a typo in the transcription of the parafermion characters in version $2$ of [BHL+21] on the arXiv and in the published version. The correct characters are $\psi^{(3)}_{1,1}(\tau)=\psi^{(3)}_{2,2}(\tau)=q^{1/30}(1+q+2q^{2}+3q^{3}+5q^{4}+7q^{5}+\dots)$ and $\psi_{2,0}^{(3)}(\tau)=q^{11/30}(1+2q+2q^{2}+4q^{3}+5q^{4}+8q^{5}+\dots)$.
The characters of $V^{\natural}$ twisted by the Fibonacci symmetry corresponding to $W$ can be computed in terms of the ordinary characters of $W$, which can in turn be computed in terms of the characters of $\mathcal{P}(3)$ and $\textsl{VF}_{24}^{\natural}$ using the identifications in equation (9). For example, using $\smash{\hat{\mathcal{L}}}$ to denote the operator acting on $\smash{V^{\natural}}$ corresponding to the non-trivial line $\mathcal{L}$ in the Fibonacci category $\mathrm{Fib}$, one obtains
@@
where $\varphi=(1+\sqrt{5})/2$. One may then compute the $\mathcal{L}$-twisted sector $\smash{V^{\natural}_{\mathcal{L}}}$ by taking a modular $S$-transformation of equation (10). Alternatively, one may proceed abstractly by noting that $\smash{V^{\natural}}$ defines a Lagrangian algebra $L$ inside $\operatorname{Rep}(W)$, and the category $\smash{\operatorname{Ver}(V^{\natural}/W)\cong\mathrm{Fib}}$ is given by $\operatorname{Rep}(W)_{L}$, the category of $L$-modules inside $\operatorname{Rep}(W)$. It is known that there are precisely two irreducible $L$-modules, $L$ and $M$, in $\operatorname{Rep}(W)$, corresponding to the trivial line and the Fibonacci line $\mathcal{L}$, respectively. They decompose as simple objects of $\operatorname{Rep}(W)$ as
@@
We mention that this vertex operator algebra $W\subset V^{\natural}$ with $\operatorname{Rep}(W)\cong Z(\mathrm{Fib})$ and character (7) has been independently discovered in unpublished work [CLY24]. There, it appears as an intermediate step in a “non-classical” construction of the moonshine module $V^{\natural}$.
Option 2. It is also possible to realize a Fibonacci symmetry of $V^{\natural}$ whose corresponding conformal subalgebra has the character vector given by (8).
Let $U_{\mathrm{2A}}$ be the unique simple-current extension of the simple Virasoro vertex operator algebras $L(\nicefrac{{1}}{{2}})\otimes L(\nicefrac{{7}}{{10}})$ (cf. 3.24) that decomposes as
$U_{\mathrm{2A}}=L(\nicefrac{{1}}{{2}})\otimes L(\nicefrac{{7}}{{10}})\oplus L(\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}})\otimes L(\nicefrac{{7}}{{10}},\nicefrac{{3}}{{2}}).$
The representation theory of this strongly rational vertex operator algebra was studied in [LY00], and one can infer from op. cit. that
$\operatorname{Rep}(U_{\mathrm{2A}})\cong(F_{4},1)\boxtimes(D_{4},1).$
As a fusion category, i.e. forgetting the braiding and twist, one can verify that $\operatorname{Rep}(U_{\mathrm{2A}})\cong\mathrm{Fib}\boxtimes\operatorname{Vect}_{\mathbb{Z}_{2}\times\mathbb{Z}_{2}}$.
It is known that $U_{\mathrm{2A}}$ embeds primitively into $V^{\natural}$, so that its commutant $\tilde{U}_{\mathrm{2A}}=\operatorname{Com}_{V^{\natural}}(U_{\mathrm{2A}})$ has representation category given by $\operatorname{Rep}(\tilde{U}_{\mathrm{2A}})\cong\overline{\operatorname{Rep}(U_{\mathrm{2A}})}$. In equation (3.55) of [BHL+21], the character vector of $\tilde{U}_{\mathrm{2A}}$ was computed, and it was argued that $\tilde{U}_{\mathrm{2A}}$ admits an action of $\!E_{6}(2).2$2 by symmetries, with the Steinberg subgroup ${{}^{2}}\!E_{6}(2)$ acting by inner automorphisms.
Now, the vertex operator algebra $U_{\mathrm{2A}}\otimes\tilde{U}_{\mathrm{2A}}$ sits as a conformal subalgebra inside $V^{\natural}$ and induces an action of $\mathrm{Ver}(V^{\natural}/U_{\mathrm{2A}}\otimes\tilde{U}_{\mathrm{2A}})\cong\mathrm{Fib}\boxtimes\operatorname{Vect}_{\mathbb{Z}_{2}\times\mathbb{Z}_{2}}$ by symmetries. [^21] It is known that each of the three non-identity involutions in the factor $\operatorname{Vect}_{\mathbb{Z}_{2}\times\mathbb{Z}_{2}}$ live in the 2A conjugacy class of $\mathrm{Aut}(V^{\natural})\cong\mathbb{M}$, where we follow the conventions of [CCN+85] in labeling conjugacy classes of the monster group. Galois theory then predicts that intermediate subalgebras sitting in between $U_{\mathrm{2A}}\otimes\tilde{U}_{\mathrm{2A}}$ and $V^{\natural}$ are in correspondence with fusion subcategories of $\mathrm{Ver}(V^{\natural}/U_{\mathrm{2A}}\otimes\tilde{U}_{\mathrm{2A}})$. In particular, one such fusion subcategory is $\mathrm{Fib}$, and we claim that the corresponding vertex operator subalgebra, $W=(V^{\natural})^{\mathrm{Fib}}$, is a simple-current extension of $U_{\mathrm{2A}}\otimes\tilde{U}_{\mathrm{2A}}$, with modules decomposing as
(11) $\begin{split}W_{(\mathbf{1},\mathbf{1})}&=U_{\mathrm{2A}}\tilde{U}_{\mathrm{2A}}\oplus U_{\mathrm{2A}}(2)\tilde{U}_{\mathrm{2A}}(2)\oplus U_{\mathrm{2A}}(3)\tilde{U}_{\mathrm{2A}}(3)\oplus U_{\mathrm{2A}}(4)\tilde{U}_{\mathrm{2A}}(4),\\ W_{(\mathbf{1},\tilde{\tau})}&=U_{\mathrm{2A}}(1)\tilde{U}_{\mathrm{2A}}\oplus U_{\mathrm{2A}}(5)\tilde{U}_{\mathrm{2A}}(2)\oplus U_{\mathrm{2A}}(6)\tilde{U}_{\mathrm{2A}}(3)\oplus U_{\mathrm{2A}}(7)\tilde{U}_{\mathrm{2A}}(4),\\ W_{(\tau,\mathbf{1})}&=U_{\mathrm{2A}}\tilde{U}_{\mathrm{2A}}(1)\oplus U_{\mathrm{2A}}(2)\tilde{U}_{\mathrm{2A}}(5)\oplus U_{\mathrm{2A}}(3)\tilde{U}_{\mathrm{2A}}(6)\oplus U_{\mathrm{2A}}(4)\tilde{U}_{\mathrm{2A}}(7),\\ W_{(\tau,\tilde{\tau})}&=U_{\mathrm{2A}}(1)\tilde{U}_{\mathrm{2A}}(1)\oplus U_{\mathrm{2A}}(5)\tilde{U}_{\mathrm{2A}}(5)\oplus U_{\mathrm{2A}}(6)\tilde{U}_{\mathrm{2A}}(6)\oplus U_{\mathrm{2A}}(7)\tilde{U}_{\mathrm{2A}}(7),\end{split}$
where we have labeled modules of $U_{\mathrm{2A}}$ and $\tilde{U}_{\mathrm{2A}}$ following equations (3.47) and (3.55) of [BHL+21]. It is straightforward to check that the graded dimensions of the modules appearing in (11) are compatible with (8) to low orders in the $q$-expansion.
@@
Let $\tilde{\alpha}=1/\sqrt{2k}$. Following [DN99b], the irreducible modules of $V^{+}_{\sqrt{k}A_{1}}$ are
$\bigl{\{}V_{\sqrt{k}A_{1}}^{\pm},V_{\sqrt{k}A_{1}+k\tilde{\alpha}}^{\pm}\bigr{\}}\cup\bigl{\{}V_{\sqrt{k}A_{1}+i\tilde{\alpha}}\bigm{|}i=1,\dots,k-1\bigr{\}}\cup\bigl{\{}V_{\sqrt{k}A_{1}}^{T_{i},\pm}\bigm{|}i=1,2\bigr{\}},$
and it is straightforward to determine that $V_{A_{1}}$ and its non-trivial irreducible module $V_{A_{1}}(\nicefrac{{1}}{{4}})$ decompose as follows. If $\ell$ is even, then
@@
$V_{A_{1}}(\nicefrac{{1}}{{4}})$ $\cong\bigoplus_{j=0}^{\ell/2-1}2\cdot V_{\ell A_{1}+(2j+1)\ell\tilde{\alpha}},$
@@
$V_{A_{1}}(\nicefrac{{1}}{{4}})$ $\cong V_{\ell A_{1}+\ell^{2}\tilde{\alpha}}^{+}\oplus V_{\ell A_{1}+\ell^{2}\tilde{\alpha}}^{-}\oplus\bigoplus_{j=0}^{(\ell-3)/2}2\cdot V_{\ell A_{1}+(2j+1)\ell\tilde{\alpha}}.$
@@
$\mathcal{H}_{\mathrm{SU}(2)_{1}}\cong V_{A_{1}}\otimes\overline{V_{A_{1}}}\oplus V_{A_{1}}(\nicefrac{{1}}{{4}})\otimes\overline{V_{A_{1}}(\nicefrac{{1}}{{4}})}$